# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v2

@@ 1.B. Organization of the paper
In Section 2 we prove Theorem A. The sublinear adaptation is mainly in Section 2.C, resulting in Proposition 2.11. The conclusion of Theorem A is in Section 2.D, and done exactly as in [KKL98].
Section 3 introduces some of the theory of completely solvable Lie groups and the significance of sublinear bilipschitz equivalence (SBE) to this theory. Corollary B and Theorem C are proved in Section 3.C. We actually deduce both of these statements from Theorem 3.16, although Corollary B could be derived directly. Theorem 3.16 is stated in terms of diagonal Heintze groups, which are the Gromov-hyperbolic groups of class $(\mathcal{C}_{1})$ and thus are of coarse type I. Finally, in Section 4 we formulate and prove Corollary 4.2 about the families of $5$-dimensional Lie groups that can be distinguished up to quasiisometry using our product theorem; corollary D appears as a byproduct of this study.
1.C. Acknowledgements
@@ 2. The product Theorem A
In this section we prove Theorem A. The argument is exactly that of Theorem B of Kapovich, Kleiner and Leeb in [KKL98]. Most of their proof works verbatim also for the case of SBE. In particular, this is true for sections 3, 4, 5 and 6 of their work. In these sections Kapovich, Kleiner and Leeb deal with properties of homeomorphisms between asymptotic cones, and to this end it makes no difference whether these homeomorphisms arise as the cone maps of quasiisometries or SBE.
The only new ingredient we need is a generalization of Section 2 of [KKL98], proving that whenever the cone maps preserve the product structure, the original maps coarsely preserve the product structure. Our Proposition 2.11 generalizes [KKL98, Proposition 2.6], and our Lemma 2.17 generalizes [KKL98, Lemma 2.7].
The key assumption in this section is that the cone maps decompose as products. A sufficient condition for this is that the original factors are of coarse type I and II [KKL98, Theorem 5.1]. This is the reason for the hypothesis of Theorem A. We only briefly sketch the definitions because in essence all we need to know about these spaces is the property that their cone maps preserve the product structure. We refer to [KKL98, Section 3] for more details.
Definition 2.1.
Definition 2.2.
Definition 2.3.
A space $X$ is of coarse type I (resp. II) if every asymptotic cone of $X$ is of type I (resp. II).
Symmetric spaces of noncompact type and higher rank are coarse type II [KL97], and so are Euclidean buildings of higher rank and cocompact affine Weyl group. Examples of spaces of coarse type I are Gromov hyperbolic groups whose Gromov boundary contains at least 3 points. These examples are enough for all our purposes in this paper, which arise in Section 3 in the proofs of Corollary B and Theorem 3.16.
We start with some preliminaries on SBE and asymptotic cones, then prove Proposition 2.11 and its variant Proposition 2.18. In Section 2.D we deduce Theorem A from these propositions and from some further results of [KKL98].
@@ 2.A. Preliminaries
We say that $u$ is admissible if it is nondecreasing and $\limsup u(2r)/u(r)<+\infty$. Often, we will also assume that $u$ is subadditive, that is, $u(r_{1}+r_{2})\leqslant u(r_{1})+u(r_{2})$ for all $r_{1},r_{2}\geqslant 0$. This implies admissibility.
@@
We say that $f$ is a $(L,u)$-bilipschitz embedding if it is $(L,cu,x_{0})$-Lipschitz and $(L,cu,x_{0})$-expansive for some $c\geqslant 0$. If $f$ is additionally $(u,y_{0})$-surjective for some $y_{0}$, then for all $y^{\prime}_{0}\in Y$ there is $c^{\prime}>0$ such that it is $(c^{\prime}u,y^{\prime}_{0})$-surjective; in this case, we say that $f$ realizes a $\big{(}L,O(u)\big{)}$-bilipschitz equivalence, or for short, a $O(u)$-bilipschitz equivalence between $X$ and $Y$. When no reference is made to $L$ and $u$ we will call a $(L,u)$-bilipschitz embedding a sublinear bilipschitz embedding.
@@ 2.B. Going through cones
Let $X$ be a metric space. Let $(\sigma_{n})_{n\in\mathbf{N}}$ be a sequence of positive real numbers. Let $(x_{n})_{n}\in X^{\mathbf{N}}$. We call precone of $X$ with data $\big{(}(x_{n})_{n},\sigma\big{)}$ the set of sequences
$\operatorname{Precone}(X,(x_{n})_{n},\sigma)=\{(x^{\prime}_{n})_{n}\in X^{\mathbf{N}}:\exists M\in[0,+\infty),\forall n\in\mathbf{N},d(x_{n},x_{n}^{\prime})\leqslant M\sigma_{n}\}.$
Given a nonprincipal ultrafilter $\omega$ over $\mathbf{N}$ we denote by $\operatorname{Cone}\big{(}X,(x_{n})_{n},(\sigma_{n})_{n},\omega\big{)}$ the quotient of the set $\operatorname{Precone}(X,(x_{n})_{n},\sigma)$ by the relation
$(x^{\prime}_{n})_{n}\sim(x^{\prime\prime}_{n})_{n}\iff\lim_{n\to\omega}\frac{d(x^{\prime}_{n},x^{\prime\prime}_{n})}{\sigma_{n}}=0$
@@
$d_{\omega}(\mathbf{[}x^{\prime}],\mathbf{[}x^{\prime\prime}])=\lim_{n\to\omega}\frac{d(x^{\prime}_{n},x^{\prime\prime}_{n})}{\sigma_{n}}.$
If $\sigma$ goes to $\infty$ we call this an asymptotic cone. When it is not relevant to mention all the remaining data, we will denote an asymptotic cone of $X$ by $X_{\omega}$.
@@
Given a metric space $X$, a triple $\big{(}X,(x_{n})_{n},(\sigma_{n})_{n}\big{)}$ is called $u$-admissible if for some (any) $v\in O(u)$
$\lim_{i\to+\infty}\frac{v(x_{i})}{\sigma_{i}}=0.$
Note that we did not specify a basepoint when writing $v(x_{i})$ in the above definition; such a choice turns out to have no influence on the notion of $u$-admissibility.
@@
Let $L\geqslant 1$. Let $u$ be an admissible function, let $X$ and $Y$ be metric spaces and let $f:X\rightarrow Y$ an $(L,u)$-bilipschitz equivalence. Assume $\big{(}X,(x_{n})_{n},(\sigma_{n})_{n}\big{)}$ is $u$-admissible. Then
$f\big{(}\operatorname{Precone}(X,x_{n},\sigma_{n})\big{)}\subseteq\operatorname{Precone}\big{(}Y,f(x_{n}),\sigma_{n}\big{)},$
@@
$\operatorname{Cone}\big{(}f,(x_{n})_{n},(\sigma_{n})_{n},\omega\big{)}\colon\operatorname{Cone}\big{(}X,(x_{n})_{n},(\sigma_{n})_{n},\omega)\to\operatorname{Cone}\Big{(}Y,\big{(}f(x_{n})\big{)}_{n},(\sigma_{n})_{n},\omega\Big{)}$
is well-defined and an $L$-bilipschitz embedding. Moreover, if $f$ is additionally $O(u)$-surjective, then $\operatorname{Cone}\big{(}f,(x_{n})_{n},(\sigma_{n})_{n},\omega)$ is a bilipschitz homeomorphism.
@@
Lemma 2.7 means that when the asymptotic cones are $u$-admissible, the cone maps of $O(u)$-bilipschitz equivalences have the same property as cone maps of quasiisometries. We therefore feel free to use results from [KKL98] whose hypothesis require some property to hold for all asymptotic cones, while we can only guarantee it for $u$-admissible cones. We do not reformulate or reprove these statements in our work, and the reader should only note that whenever we use cones, they are $u$-admissible.
@@
Let $X,Y$ be two metric spaces, $v:\mathbf{R}_{\geqslant 0}\rightarrow\mathbf{R}_{\geqslant 1}$. Two maps $f_{1},f_{2}:X\rightarrow Y$ are at distance $v$ if for all $x\in X$ one has $d_{Y}\big{(}f_{1}(x),f_{2}(x)\big{)}\leq v(|x|)$. The maps are called sublinearly close if $f_{1},f_{2}$ are at distance $v$ for some sublinear function $v$.
2.C. SBE Preserve Product Structure
Definition 2.10.
A bilipschitz homeomorphism $F:\prod_{i=1}^{n}X_{i}\rightarrow\prod_{i=1}^{m}Y_{i}$ is said to preserve the product structure if up to reindexing the factors and ignoring factors not in the image of $F$, there are bilipschitz homeomorphisms $F_{i}$ such that for each $i$, the following diagram commutes:
$\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{F}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{F_{i}}$
Throughout this section, the standing assumptions on the map $\phi$ and spaces $X,Y$ are those of Proposition 2.11.
Proposition 2.11 (SBE version of Proposition 2.6 in [KKL98]).
Suppose $\phi:X=\prod_{i}X_{i}\rightarrow Y=\prod Y_{j}$ is an $(L,u)$-bilipschitz equivalence. Assume that for all $u$-admissible cones $X_{\omega}$ and $Y_{\omega}$ of $X$ and $Y$, the ultralimit $F=\operatorname{Cone}(\phi):X_{\omega}\rightarrow Y_{\omega}$ preserves the product structure. Then:
  1. (1)
    Up to reindexing the factors, there are $(L^{\prime},u_{i})$-bilipschitz equivalences $\phi_{i}:X_{i}\rightarrow Y_{i}$ with $u_{i}\in O(u)$.
  2. (2)
    The diagram
    $\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\phi_{i}}$
    commutes up to $O(u)$ sublinear error. In particular, $\phi$ is sublinearly close to a product of $O(u)$-bilipschitz equivalences.
Remark 2.12.
One notable difference between our proofs and those of the product theorem for quasiisometries concerns the ‘nontranslatability’ condition [KKL98, Definitions 2.2, 2.3]. In Section 2 of [KKL98] the nontranslatability of the factors was used both implicitly in the assumption that cone maps preserve product structure, and then again explicitly in order to prove that certain pairs of quasi-isometries are at uniform bounded distance. In contrast, we use the nontranslatability only through its implicit use as part of the product decomposition of cone maps between spaces of types I and II. We use this result as a black box from [KKL98] and therefore do not need to explicitly define or discuss nontranslatability. In particular our hypothesis in Proposition 2.11 is formally (and perhaps essentially) weaker than its quasiisometry predecessor [KKL98, Proposition 2.6].
The proof requires some sublinear adaptations to the notions that appear in [KKL98].
Definition 2.13.
Let $d>0$ be a positive constant and $d_{0}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 1}$ a function. A pair of points $x,x^{\prime}\in X$ is called $d$-separated if $d_{X}(x,x^{\prime})\geqslant d$, and $d_{0}$-separated if $d(x,x^{\prime})\geqslant d_{0}(|x|\vee|x^{\prime}|)$.
@@ 2.C. SBE Preserve Product Structure
Fix $L_{1}>L$, $\epsilon<L_{1}^{-1}$, $i,j\in\{1,\dots,n\}$. A pair of $i$-horizontal points $x,x^{\prime}\in X$ is called:
@@
    $j$-compressed if
    $\frac{d_{Y_{j}}\Big{(}\pi_{Y_{j}}\big{(}\phi(x)\big{)},\pi_{Y_{j}}\big{(}\phi(x^{\prime})\big{)}\Big{)}}{d_{X}(x,x^{\prime})}\leq\epsilon$
@@
    $j$-uncompressed if
    $L_{1}^{-1}\leq\frac{d_{Y_{j}}\Big{(}\pi_{Y_{j}}\big{(}\phi(x)\big{)},\pi_{Y_{j}}\big{(}\phi(x^{\prime})\big{)}\Big{)}}{d_{X}(x,x^{\prime})}\leq L_{1}$
@@
    $j$-semi-compressed if
    $\epsilon\leq\frac{d_{Y_{j}}\Big{(}\pi_{Y_{j}}\big{(}\phi(x)\big{)},\pi_{Y_{j}}\big{(}\phi(x^{\prime})\big{)}\Big{)}}{d_{X}(x,x^{\prime})}\leq{L_{1}}^{-1}$
@@
When it is clear from the context, we occasionally do not specify the space in which the distance is computed. Also, we sometimes write $\pi_{j}$ instead of $\pi_{X_{j}}$ or $\pi_{Y_{j}}$.
Let $x,x^{\prime}$ be $i$-horizontal. Notice two trivial facts: a) $d_{X_{i}}\big{(}\pi_{i}(x),\pi_{i}(x^{\prime})\big{)}=d_{X}\big{(}x,x^{\prime})$, and b) the geodesic $x,x^{\prime}]$ in $X$ joining $x$ and $x^{\prime}$ is contained in a leaf of $X_{i}$. In particular, for each $x^{\prime\prime}\in[x,x^{\prime}]$, $(x,x^{\prime\prime})$ and $(x^{\prime},x^{\prime\prime})$ are $i$-horizontal pairs.
@@
Lemma 2.17 (SBE version of Lemma $2.7$ in [KKL98]).
There exists a sublinear subadditive function $d_{0}\in O(u)$ such that for every fixed $i,j\in\{1,\dots,n\}$, either all $d_{0}$-separated $i$-horizontal pairs are $j$-compressed or all such pairs are $j$-uncompressed.
The existence of $d_{0}$ as in Lemma 2.17 in particular implies that $d_{0}$-separated $i$-horizontal pairs are never $j$-semi-compressed. Our proof follows that of Lemma 2.7 in [KKL98], with the required sublinear adaptations. The meaning of the lemma is that $\phi$ coarsely preserve the product structure of $X$ and $Y$. The idea of the proof is that if $\phi$ fails to do so, this failure is manifested by a sequence of points and constants that give rise to $u$-admissible asymptotic cones of $X$ and $Y$ in which $\operatorname{Cone}(\phi)$ also fails to preserve the product structure, contradicting the hypothesis.

### v3

@@ 1.B. Organization of the paper
The proofs of Theorems A and C are in Section 2. Section 3 introduces some of the theory of completely solvable Lie groups and the significance of sublinear bilipschitz equivalence to this theory. Corollary B and Theorem D are proved in Section 3.C. We actually deduce both of these statements from Theorem 3.16, although Corollary B could be derived directly. Theorem 3.16 is stated in terms of diagonal Heintze groups, which are the Gromov-hyperbolic groups of class $(\mathcal{C}_{1})$ and thus are of coarse type I. Finally, in Section 4 we formulate and prove Corollary 4.2 about the families of $5$-dimensional Lie groups that can be distinguished up to quasiisometry using our product theorem; Corollary E appears as a byproduct of this study, in Corollary 4.4.
1.C. Acknowledgments
@@ 2. The product Theorem A
In this section we prove Theorem C and conclude Theorem A. We start with some preliminaries on sublinear bilipschitz equivalence and sublinear functions. Then we present asymptotic cones, which are the key objects in the proof. We then turn to prove Theorem C and formulate (without proof) a variant which allows us to deal with Euclidean factors in Proposition 2.22. In Section 2.D we deduce Theorem A. The strategy appears already in [KKL98], and most of our work is in adapting it to the sublinear setting.
@@ 2.A. Preliminaries
We say that $u$ is admissible if it is nondecreasing and $\limsup u(2r)/u(r)<+\infty$. Often, we will also assume that $u$ is subadditive, that is, $u(r_{1}+r_{2})\leqslant u(r_{1})+u(r_{2})$ for all $r_{1},r_{2}\geqslant 0$. This implies admissibility. This assumption is not extremely restrictive, as we prove in Lemma 2.2 below: a sublinear nondecreasing function is always dominated by a subadditive (in particular, admissible) one. We first need an intermediate technical statement.
Lemma 2.1.
Let $u_{0}\colon\mathbf{R}_{\geqslant 0}\to\mathbf{R}_{\geqslant 0}$ be such that $\lim_{r\to+\infty}\frac{u_{0}(r)}{r}=0$, $u_{0}(0)=0$, and $\limsup_{r\to 0}\frac{u_{0}(r)}{r}<+\infty$. Then there exists $v_{0}\colon\mathbf{R}_{\geqslant 0}\to\mathbf{R}_{\geqslant 0}$ such that $u_{0}\leqslant v_{0}$ and $v_{0}$ is a concave function such that $\lim_{r\to+\infty}\frac{v(r)}{r}=0$.
Proof.
Let $\mathcal{W}$ be the set of concave functions $w$ from $\mathbf{R}_{\geqslant 0}$ to itself and such that $u_{0}\leqslant w$. Given the assumptions that we made on $u_{0}$, $\mathcal{W}$ contains a linear function, especially it is not empty. For all $r\in\mathbf{R}_{\geqslant 0}$ we set $v_{0}(r)=\inf\{w(r)\colon w\in\mathcal{W}\}$ and proceed to prove that the function $v_{0}$ is sublinear. If it was not the case, there would be $\varepsilon>0$ and a sequence $\{r_{i}\}_{i\in\mathbf{N}}$ going to infinity, such that $v_{0}(r_{i})\geqslant\varepsilon r_{i}$. Note however that $v_{0}$ is concave, as it is an infimum of concave functions. So from the previous inequality and the convexity inequality, we get that $v_{0}(r)\geqslant r$ for every nonnegative $r$. But $u_{0}$ is sublinear, so there exists $R>0$ such that $u_{0}(r)\leqslant\varepsilon r/2$ for every $r\geqslant R$. Considering the function $v_{1}$ such that $v_{1}(r)=\epsilon r$ if $0\leqslant r\leqslant R$ and $v_{1}(r)=\varepsilon R/2+\varepsilon r/2$ if $r\geqslant R$, we get a function $v_{1}$ in $\mathcal{W}$ eventually smaller than $v_{0}$, a contradiction. This terminates the proof that $v_{0}$ is sublinear; the other required properties of $v_{0}$ have been established in the course of the proof. ∎
Lemma 2.2 (Dominating sublinear increasing functions by subadditive ones).
Let $u\colon\mathbf{R}_{\geqslant 0}\to\mathbf{R}_{\geqslant 1}$ be a nondecreasing sublinear function. There exists a subadditive, nondecreasing, sublinear function $v$ such that $u\leqslant v$.
Proof.
Given the assumptions made on $u$, there exists $M\geqslant 0$ such that $u-M\leqslant u_{0}$, where $u_{0}$ is a sublinear function such that $u_{0}(0)=0$ and $\limsup_{r\to 0}\frac{u_{0}(r)}{r}=0$. Applying Lemma 2.1 to $u_{0}$ we get a concave function $v_{0}$ dominating $u_{0}$. Now set $v_{1}=v_{0}+M$. We claim that $v_{1}$ is subadditive. Indeed, $v_{0}$ is concave and $v_{0}(0)=0$, so that for every pair $(r,s)$ of nonnegative real numbers, not both equal to zero, we have
$v_{0}(r)\geqslant\frac{r}{r+s}v_{0}(r+s)\;\text{and}\;v_{0}(s)\geqslant\frac{s}{r+s}v_{0}(r+s)$
which after summation gives that $v_{0}$ is subadditive; now, for any $r,s\geqslant 0$,
$v_{1}(r+s)$ $=v_{0}(r+s)+M$
$\leqslant v_{0}(r)+v_{0}(s)+M$
$\leqslant v_{0}(r)+M+v_{0}(s)+M$
$=v_{1}(r)+v_{1}(s).$
In addition, $v_{0}$ is sublinear, so $v_{1}$ is as well. Finally, assume towards contradiction that there exists $(r,s)$ such that $0\leqslant r<s$ and $_{1}(r)>v_{1}(s). vv$ being concave, it will then eventually take negative values, as $v_{1}(t)\leqslant v_{1}(r)-\frac{t-r}{s-r}(v_{1}(r)-v_{1}(s))$ which becomes negative for large enough $t$. This is a contradiction, hence $v_{1}$ is nondecreasing. One may take $v=v_{1}$ and the conclusion is reached. ∎
Remark 2.3.
We do not have $v\in O(u)$ in the conclusion of lemma 2.2, as the following example shows: $u(t)=\exp(\lfloor\log(\log(t))\rfloor^{2})$ for large enough $t$. Then for $t_{n}=e^{e^{n}}-1/2$ we have $\lim_{n}u(t_{n}+1)/u(t_{n})=+\infty$, and $\lim_{n}v(t_{n}+1)/v(t_{n})=+\infty$ for any $v\in O(u)$ dominating $u$, which cannot be for a subadditive $v$.
Remark 2.4.
The class of admissible functions is stable under the operations $\vee$ and $\wedge$. This will be of use in Section 3 where we will use that if $u$ is admissible, then $u\vee\log$ is admissible.
@@
We say that $f$ is an $(L,u)$-bilipschitz embedding if it is $(L,cu,x_{0})$-Lipschitz and $(L,cu,x_{0})$-expansive for some $c\geqslant 0$. If $f$ is additionally $(u,y_{0})$-surjective for some $y_{0}$, then for all $y^{\prime}_{0}\in Y$ there is $c^{\prime}>0$ such that it is $(c^{\prime}u,y^{\prime}_{0})$-surjective; in this case, we say that $f$ realizes an $\big(L,O(u)\big)$-bilipschitz equivalence, or for short, an $O(u)$-bilipschitz equivalence between $X$ and $Y$. In line with the terminology for quasiisometries, $L$ is sometimes called the bilipschitz constant of $f$. When no reference is made to $L$ and $u$ we will call an $(L,u)$-bilipschitz embedding a sublinear bilipschitz embedding. We may omit the base points or any of the constants when they are clear from the context.
@@ 2.B. Going through cones
Let $X$ be a metric space. Let $(\sigma_{k})_{k\in\mathbf{N}}$ be a sequence of positive real numbers. Let $(x_{k})_{k}\in X^{\mathbf{N}}$. We call precone of $X$ with data $\big((x_{k})_{k},(\sigma_{k})_{k}\big)$ the set of sequences
$\operatorname{Precone}\big(X,(x_{k})_{k},(\sigma_{k})_{k}\big)=\{(x^{\prime}_{k})_{k}\in X^{\mathbf{N}}:\exists M\in[0,+\infty),\forall k\in\mathbf{N},d(x_{k},x_{k}^{\prime})\leqslant M\sigma_{k}\}.$
Given a nonprincipal ultrafilter $\omega$ over $\mathbf{N}$ we denote by $\operatorname{Cone}\big(X,(x_{k})_{k},(\sigma_{k})_{k},\omega\big)$ the quotient of the set $\operatorname{Precone}\big(X,(x_{k})_{k},(\sigma_{k})_{k}\big)$ by the relation
$(x^{\prime}_{k})_{k}\sim(x^{\prime\prime}_{k})_{k}\iff\lim_{\omega}\frac{d(x^{\prime}_{k},x^{\prime\prime}_{k})}{\sigma_{k}}=0$
@@
$d_{\omega}(\mathbf{[}x^{\prime}],\mathbf{[}x^{\prime\prime}])=\lim_{\omega}\frac{d(x^{\prime}_{k},x^{\prime\prime}_{k})}{\sigma_{k}}.$
If $\sigma_{k}$ goes to $\infty$ we call this an asymptotic cone.
When it is not relevant to mention all the remaining data, we will denote an asymptotic cone of $X$ by $X_{\omega}$. Often the defining data is relevant, in which case we ease notation and write $\operatorname{Cone}(X,x_{k},\sigma_{k})$.
@@
Given a metric space $X$, a triple $(X,x_{k},\sigma_{k})$ is called $u$-admissible if for some (any) $v\in O(u)$
$\lim_{k\to+\infty}\frac{v(|x_{k}|)}{\sigma_{k}}=0.$
An asymptotic cone $X_{\omega}:=\operatorname{Cone}(X,x_{k},\sigma_{k})$ is called $u$-admissible if the triple $(X,x_{k},\sigma_{k})$ is $u$-admissible.
Note that we did not specify a basepoint when writing $v(|x_{k}|)$ in the above definition; such a choice turns out to have no influence on the notion of $u$-admissibility.
@@
Let $L\geqslant 1$. Let $u$ be an admissible function, let $X$ and $Y$ be metric spaces and let $f:X\rightarrow Y$ an $(L,u)$-bilipschitz equivalence. Assume $\big(X,(x_{k})_{k},(\sigma_{k})_{k}\big)$ is $u$-admissible. Then
$f\Big(\operatorname{Precone}\big(X,(x_{k})_{k},(\sigma_{k})_{k}\big)\Big)\subseteq\operatorname{Precone}\Big(Y,\big(f(x_{k})\big)_{k},(\sigma_{k})_{k}\Big),$
@@
$\operatorname{Cone}\big(f,(x_{k})_{k},(\sigma_{k})_{k},\omega\big)\colon\operatorname{Cone}\big(X,(x_{k})_{k},(\sigma_{k})_{k},\omega)\to\operatorname{Cone}\Big(Y,\big(f(x_{k})\big)_{k},(\sigma_{k})_{k},\omega\Big)$
is well-defined and an $L$-bilipschitz embedding. Moreover, if $f$ is additionally $O(u)$-surjective, then $\operatorname{Cone}\big(f,(x_{k})_{k},(\sigma_{k})_{k}\big)$ is a bilipschitz homeomorphism.
@@
The notion of $u$-admissibility concerns a key difference between quasiisometries and $O(u)$-bilipschitz equivalence. When $u\in O(1)$, then every asymptotic cone is $u$-admissible, and in particular the cone maps of quasiisometries are bilipschitz homeomorphisms in every asymptotic cone. However when $u\notin O(1)$, not all cone maps are bilipschitz homeomorphisms. In our work, this adds a significant difficulty in the context of controlling the error term in Theorem C (see Remarks 1.6 and 1.7).
Remark 2.10.
Unless stated otherwise, whenever we define $X_{\omega}=\operatorname{Cone}(X,x_{k},\sigma_{k})$, we implicitly define $Y_{\omega}:=\operatorname{Cone}\big(Y,f(x_{k}),\sigma_{k}\big)$. Lemma 2.8 implies that if $X_{\omega}$ is $u$-admissible, so is $Y_{\omega}$.
2.C. Proof of Theorem C
Throughout this section, the standing assumptions on the map $\phi$ and spaces $X,Y$ are those of Theorem C. When it is clear from the context, we occasionally do not specify the space in which the distance is computed. Also, we sometimes write $\pi_{j}$ instead of $\pi_{X_{j}}$ or $\pi_{Y_{j}}$.
The following are some sublinear adaptations to the notions that appear in [KKL98].
@@ 2.C. Proof of Theorem C
Let $d>0$ be a positive constant and $d_{0}:\mathbf{R}_{\geq 0}\rightarrow\mathbf{R}_{\geq 1}$ a function. A pair of points $x,x^{\prime}\in X$ is called $d$-separated if $d_{X}(x,x^{\prime})\geqslant d$, and $d_{0}$-separated if $d(x,x^{\prime})\geqslant d_{0}(|x|\vee|x^{\prime}|)$.
@@
Let $x,x^{\prime}\in X$ be a pair of $i$-horizontal points, and fix $L_{1}>1,\epsilon<L_{1}^{-1}$. The following definitions depend on $L_{1}$ and $\epsilon$, but we suppress them from the terminology:
@@
    $(x,x^{\prime})$ is $j$-compressed if
    dYj(πYj(ϕ(x)),πYj(ϕ(x’)))dX(x,x’)≤ϵ
@@
    $(x,x^{\prime})$ is $j$-uncompressed if
    L_1^-1≤dYj(πYj(ϕ(x)),πYj(ϕ(x’)))dX(x,x’)≤L_1
@@
    $(x,x^{\prime})$ is $j$-semi-compressed if ϵ≤dYj(πYj(ϕ(x)),πYj(ϕ(x’)))dX(x,x’)≤L_1^-1.
@@
Recall that $L$ is the bilipschitz constant of the map $\phi:X\rightarrow Y$. We fix once and for all $L_{1}>L$ and $\epsilon<L_{1}^{-1}$, and use the terminology of Definition 2.13 with respect to these constants. Our proof of Theorem C works for any choice of $L_{1}>L$ and $\epsilon<L_{1}^{-1}$ (compare below with Remark 2.23).
For an $i$-horizontal pair $x,x^{\prime}\in X$, we note two trivial facts:
  1. (1)
    $d_{X_{i}}\big(\pi_{i}(x),\pi_{i}(x^{\prime})\big)=d_{X}\big(x,x^{\prime})$.
  2. (2)
    Any geodesic $x,x^{\prime}]$ in $X$ joining $x$ and $x^{\prime}$ is contained in a fiber of $X_{i}$. In particular, for each $x^{\prime\prime}\in[x,x^{\prime}]$, $(x,x^{\prime\prime})$ and $(x^{\prime},x^{\prime\prime})$ are $i$-horizontal pairs.
@@
The following lemma is straightforward and will be used repeatedly.
Lemma 2.15.
Let $X=\prod_{i=1}^{n}X_{i},Y=\prod_{i=1}^{n}Y_{i}$, $\phi:X\rightarrow Y$ an $(L,u)$-bilipschitz equivalence. Let $X_{\omega}=\operatorname{Cone}(X,x_{k},\sigma_{k}),Y_{\omega}=\operatorname{Cone}(Y,\phi(x_{k}),\sigma_{k})$ be $u$-admissible asymptotic cones, and assume $\phi_{\omega}=\operatorname{Cone}(\phi):X_{\omega}\rightarrow Y_{\omega}$ preserves the product structure, with factor maps $\phi_{\omega_{i}}:X_{\omega_{i}}\rightarrow Y_{\omega_{i}}$. Let $z_{\omega}:=[(z_{k})_{k}],z_{\omega}^{\prime}:=[(z_{k}^{\prime})_{k}]\in X_{\omega}$. It holds that
  1. (1)
    If $\lim_{\omega}\frac{1}{\sigma_{k}}d_{X_{j}}\big(\pi_{j}(z_{k}),\pi_{j}(z_{k}^{\prime})\big)=0$ then $\lim_{\omega}\frac{1}{\sigma_{k}}d_{Y_{j}}\Big(\pi_{j}\big(\phi(z_{k})\big),\pi_{j}\big(\phi(z_{k}^{\prime})\big)\Big)=0$.
  2. (2)
    If for all $l\neq i$ $\lim_{\omega}\frac{1}{\sigma_{k}}d_{X_{l}}\big(\pi_{l}(z_{k}),\pi_{l}(z_{k}^{\prime})\big)=0$, then limdYi(πi(ϕ(zk))i(ϕ(zk’)))dX(zk,zk’)[ 1L,L ] .
In particular, if the pairs $(z_{k},z_{k}^{\prime})$ which define $z_{\omega}$ and $z_{\omega}^{\prime}$ are $\omega$-almost surely $i$-horizontal, then $\omega$-almost surely $(z_{k},z_{k}^{\prime})$ is $j$-compressed for all $j\neq i$ and $i$-uncompressed.